# Denoising diffusion probabilistic models are optimally adaptive to unknown low dimensionality

### v2

@@ 4.1.1 Low dimensionality of the target distribution
Given that we focus on data residing in the Euclidean space, we can measure the complexity through the Euclidean metric entropy below (Wainwright,, 2019, Chapter 5):
@@
Following Li and Yan, 2024a , we make the following assumption to characterize the complexity of the data distribution. We shall often refer to $k$ as the intrinsic dimension of $\mathcal{X}_{\mathsf{data}}$.
@@
$\sup_{x\in\mathcal{X}_{\mathsf{data}}}\|x\|_{2}\leq R,\qquad\mathrm{where}\enspace R\coloneq k^{C_{R}}.$
@@
  • Linear subspace. The metric entropy of a bounded subset of any $k$-dimensional linear subspace is at most $O(k\log(1/\varepsilon_{0}))$ (Vershynin,, 2018, Section 4.2.1), and hence Assumption 1 covers linear subspaces studied by some prior work (e.g., Chen et al., 2023b ).
  • Low-dimensional manifold. Assumption 1 covers various manifolds under mild regularity conditions, which have been commonly used in prior work to model low-dimensional structure. To make it precise, let us first introduce some geometric quantities of the manifold, whose geometric interpretations can be found in Lee, (2018, Chapter 6). Suppose $\mathcal{M}\subseteq\mathbb{R}^{d}$ to be an isometrically embedded, compact, $k$-dimensional submanifold, then we can define:
@@
    Roughly speaking, the reach of a manifold measures the scale of which the manifold can be regarded as a linear subspace. The larger the reach $\tau_{\mathcal{M}}$ is, the more regular the manifold appears. The volume of the manifold $\mathcal{M}$ is a measure of the size of the manifold, which generalizes the concept of volume in 3-dimensional Euclidean space. Instead of requiring data to be strictly supported on the manifold, Assumption 1 also allows the data to lie in the $\varepsilon_{0}$-enlargement of the manifold $\mathcal{M}$ defined as $\mathcal{M}^{\varepsilon_{0}}:=\{x\in\mathbb{R}^{d}\mid\mathsf{dist}(x,\mathcal{M})\leq\varepsilon_{0}\}$. This is stated formally in the following lemma, with the proof deferred to Appendix C.1.
@@
  • A general set with doubling dimension $k$. Another notion for intrinsic dimension used in statistics for a general set in $\mathbb{R}^{d}$ is the doubling dimension (Dasgupta and Freund,, 2008), as defined below.
@@
    This notion of dimension can also be incorporated into our framework, as demonstrated by the following lemma (Kpotufe and Dasgupta,, 2012, Lemma 6):
@@
  • Union of the above sets. Our framework can also cover a union of the structures mentioned above, that is, Assumption 1 automatically holds when the data are supported on the union of $k^{O(k)}$ subsets, each of which satisfies Assumption 1.
@@ 4.1.2 Discretization time points
Recall that the discretization time points are $0=t_{0}<t_{1}<\cdots<t_{N}<t_{N+1}=T$, with $T$ the time horizon of the forward process. In addition, we also define throughout the early stopping time
@@
Following prior diffusion model theory (e.g., Benton et al., (2024)), we introduce a key quantity, $\kappa$, defined to be the smallest quantity obeying
@@
To get a sense of the size of $\kappa$, it is helpful to bear in mind the following schedule that has been adopted in recent theoretical analysis (Chen et al., 2023a, ; Benton et al.,, 2024; Li et al., 2024b, ):
$t_{n}$ =$=\begin{cases}\frac{2(T-1)n}{N},\qquad&\text{ if }0\leq n\leq N/2,\\ T-\delta^{2(n-N/2)/N},&~{}\text{if }N/2<n\leq N,\end{cases}$ (16)
where we assume $N$ to be even for simplicity. In words, this schedule of time points can be divided into two phases: the first phase consists of the first $N/2$ time instances that are linearly spaced within $[0,T-1]$, whereas the second phase consists of the remaining $N/2$ time points that whose distance to $T$ decays exponentially from 1 to $\delta$. As shown in Benton et al., (2024, Corollary 1), this choice (16) satisfies
@@
For this reason, we can often interpret $\kappa$ as the average time interval, or simply $\widetilde{O}(1/N)$ as $T$ is typically taken to be $\widetilde{O}(1)$.
@@ 4.1.3 Score estimation errors
Assume that the weighted sum of the mean squared score estimation errors is finite, i.e.,
@@
Treating the score learning phase as a black box, Assumption 3 singles out a single quantity, $\varepsilon_{\mathsf{score}}\geq 0$, that captures the accuracy of the pre-trained score functions, which is commonly used in recent development of diffusion model theory (Chen et al., 2023a, ; Li et al.,, 2023; Li et al., 2024c, ; Benton et al.,, 2024; Dou et al.,, 2024; Li and Yan, 2024b, ; Li et al., 2024a, ; Li and Yan, 2024a, ; Potaptchik et al.,, 2024; Li and Jiao,, 2024; Azangulov et al.,, 2024; Liang et al.,, 2024). Notably, assuming a weighted average of the mean squared score errors is clearly less stringent than requiring the mean squared score estimation errors to hold uniformly over all time discretization points (Chen et al., 2022b, ; Lee et al.,, 2022; Chen et al., 2023c, ; Wu et al.,, 2024).
@@ 4.2 Convergence guarantees for DDPM
  • $p_{\mathsf{output}}$: the distribution of $Y_{t_{N}}$ (or equivalently, the distribution of the DDPM output after $N$ steps).
@@
Suppose that Assumptions 1, 2 and 3 hold. Assume that $T>1$, $\kappa\leq 0.9$, $11/\mathrm{poly}(k)\lesssim\delta<$ 1 and $d=O\big{(}\mathrm{poly}(k)\big{)}$. Then the DDPM sampler (2) satisfies
$\mathsf{KL}(q_{\delta}\,||\,p_{\mathsf{output}})\lesssim\kappa\min\left\{\mathbb{E}\big{[}\|X_{0}\|_{2}^{2}\big{]},Tk\log k\right\}+\kappa^{2}Nk\log k+\varepsilon_{\mathsf{score}}^{2}+\big{(}d+\mathbb{E}\big{[}\|X_{0}\|_{2}^{2}\big{]}\big{)}e^{-2T}.$ (18)
The KL divergence upper bound in Theorem 2 can be divided into three parts. The first two terms on the right-hand side of (18) capture the time discretization error of the DDPM dynamics (namely, the error incurred when discretizing (13) as (11b)), the third term on the right-hand side of (18) reflects the effect of the score estimation error, whereas the fourth term corresponds to the initialization error (namely, the error induced by drawing $Y_{0}$ from $\mathcal{N}(0,I_{d})$ rather than the distribution of $X_{T}$). Note that we are comparing the DDPM output with $q_{\delta}$ (i.e., the distribution of a slightly noisy version of the data) rather than $p_{\mathsf{data}}=q_{0}$, as adding a little Gaussian noise helps avoid some highly irregular cases. For this reason, $\delta>0$ is commonly referred to as the early stopping time in prior theory. Note that $\delta$ can be taken to be very small, ensuring that $q_{\delta}$ is exceedingly close to $p_{\mathsf{data}}$.
@@
$\mathsf{KL}(q_{\delta}\,||\,p_{\mathsf{output}})\lesssim\varepsilon^{2}+\varepsilon_{\mathsf{score}}^{2},$
provided that $11/\mathrm{poly}(k)\lesssim\delta<$ 1, $d=O\big{(}\mathrm{poly}(k)\big{)}$, and
@@
$N=\widetilde{\Theta}(k/\varepsilon^{2})~{}\text{steps}$
@@
to approximate the distribution of $X_{\delta}$ (i.e., the target data distribution $p_{\mathsf{data}}$ convoluted with exceedingly small Gaussian noise) to within $\varepsilon^{2}$ KL divergence. Encouragingly, this iteration complexity is linear in the intrinsic dimension $k$ (modulo some logarithmic factor), which reduces to the state-of-the-art theory when $k=d$ (i.e., the case without low-dimensional structure) (Benton et al.,, 2024). It is noteworthy that all this is achieved without any prior knowledge about the potential low-dimensional structure in $p_{\mathsf{data}}$, thereby unraveling the remarkable capability of DDPM to adapt to unknown low dimensionality. It is also worth noting that this linear dependency in $k$ holds without any additional burn-in requirement; that is, its validity is guaranteed without requiring $N$ to first exceed, say, some polynomial function in $k$ or $d$.
@@
To understand the optimality of DDPM in adapting to low dimensionality, it is helpful to first revisit the lower bounds established by Chen et al., 2022b (, Theorem 7) and Benton et al., (2024, Appendix H) for the full-dimensional case $k=d$; namely, there exists a simple $d$-dimensional distribution $p_{\mathsf{data}}$ for which the KL error is lower bounded by some function scaling linearly in $d$. These arguments can be straightforwardly extended to the low-dimensional case, revealing the existence of a distribution $p_{\mathsf{data}}$ with intrinsic dimension $k$, that requires an iteration complexity scaling at least linearly with $k$, regardless of the sampler in use.
Comparison with the concurrent work Potaptchik et al., (2024).
@@
Assuming the data lies on a low-dimensional manifold, this concurrent work, posted two weeks before ours, established an iteration complexity that scales linearly in the manifold dimension. As alluded to previously, our results and theirs share clearly similarities in both the aims and the results. Here, let us also point out several differences here. In Potaptchik et al., (2024), the data is assumed to be strictly supported on a constant-scale manifold with diameter less than $1$ (see Assumptions A and B therein), and the density function (w.r.t. to the standard volume measure) needs to be both upper and lower bounded uniformly (see Assumption C therein). In comparison, we consider a more general setting where the data support is only required to stay close to a low-dimensional structure whose size can grow polynomially in $k$. Also, no boundedness requirement is imposed on the density function in our theory, thereby accommodating a much broader scenario. Notably, the theory in Potaptchik et al., (2024) only applies to data with $\mathbb{E}\big{[}\|X_{0}\|_{2}^{2}\big{]}=O(1)$ second moment (as a consequence of their assumption that the diameter of the manifold is $O(1)$), while ours is able to accommodating polynomia lly large second moment $\mathbb{E}\big{[}\|X_{0}\|_{2}^{2}\big{]}=\mathrm{poly}(k)$. Our proof invokes the covering-style argument to accommodate broader low-dimensional structure, which differs from the techniques in Potaptchik et al., (2024).

### v3

@@ 4.1.1 Low dimensionality of the target distribution
Given that we focus on data residing in the Euclidean space, we can measure the complexity through the Euclidean metric entropy below defined in Wainwright (2019, Chapter 5):
@@
Following Li and Yan (2024a), we make the following assumption to characterize the complexity of the data distribution, which generalizes the data generation assumption in Li and Yan (2024a). We shall often refer to $k$ as the intrinsic dimension of $\mathcal{X}_{\mathsf{data}}$.
@@
$\sup_{x\in\mathcal{X}_{\mathsf{data}}}\|x\|_{2}\leq R,\qquad\mathrm{where}\kern 5.0ptR\coloneq k^{C_{R}}.$
@@
  • Linear subspace. The metric entropy of a bounded subset of any $k$-dimensional linear subspace is at most $O(k\log(1/\varepsilon_{0}))$ (Vershynin, 2018, Section 4.2.1), and hence Assumption 1 covers linear subspaces studied by some prior work (e.g., Chen et al. (2023b)).
  • Low-dimensional manifold. Assumption 1 covers various manifolds under mild regularity conditions, which have been commonly used in prior work to model low-dimensional structure. To make it precise, let us first introduce some geometric quantities of the manifold, whose geometric interpretations can be found in Lee (2018, Chapter 6). Suppose $\mathcal{M}\subseteq\mathbb{R}^{d}$ to be an isometrically embedded, compact, $k$-dimensional submanifold, then we can define:
@@
    Figure 1: Illustration of reach $\tau_{\mathcal{M}}$ of a manifold $\mathcal{M}$ embedded in $\mathbb{R}^{2}$.
    Roughly speaking, the reach of a manifold measures the scale of which the manifold can be regarded as a linear subspace. An illustration of the reach is provided in Figure 1. The larger the reach $\tau_{\mathcal{M}}$ is, the more regular the manifold appears. The volume of the manifold $\mathcal{M}$ is a measure of the size of the manifold, which generalizes the concept of volume in 3-dimensional Euclidean space. Instead of requiring data to be strictly supported on the manifold, Assumption 1 also allows the data to lie in the $\varepsilon_{0}$-enlargement of the manifold $\mathcal{M}$ defined as $\mathcal{M}^{\varepsilon_{0}}:=\{x\in\mathbb{R}^{d}\mid\mathsf{dist}(x,\mathcal{M})\leq\varepsilon_{0}\}$. This is stated formally in the following lemma, with the proof deferred to Appendix C.1.
@@
  • A general set with doubling dimension $k$. Another notion for intrinsic dimension used in statistics for a general set in $\mathbb{R}^{d}$ is the doubling dimension (Dasgupta and Freund, 2008), as defined below.
@@
    This notion of dimension can also be incorporated into our framework, as demonstrated by the following lemma (Kpotufe and Dasgupta, 2012, Lemma 6):
@@
  • A general set with Minkowski dimension $k$. As a commonly used concept of intrinsic dimension, the Minkowski dimension is defined as follows (Xia and Vershynina, 2010).
    Definition 4 (Minkowski dimension).
    The Minkowski dimension of $\mathcal{X}_{\mathsf{data}}$ can be defined as
    $\lim_{\varepsilon\to 0^{+}}-\frac{\log N^{\mathsf{cover}}(\mathcal{X}_{\mathsf{data}},\|\cdot\|_{2},\varepsilon)}{\log\varepsilon},$
    provided that the above limit exists.
    Interestingly, our Assumption 1 can be regarded as a weaker version of the Minkowski dimension, as stated in the following lemma.
    Lemma 3.
    If $\mathcal{X}_{\mathsf{data}}$ is contained within an $\varepsilon_{0}$-enlargement of a set that has Minkowski dimension $k$, then Assumption 1 holds.
    Remark 3.
    In some literature (Falconer et al., 2023), the Minkowski dimension is defined through the limit scale of the box-covering number (denoted by $N^{\mathsf{box}\text{-}\mathsf{cover}}$) or the packing number (denoted by $N^{\mathsf{pack}}$), instead of the ball-covering number (denoted by $N^{\mathsf{cover}}$) used here. We remark here that all three definitions of the Minkowski dimension are equivalent, and Lemma 3 holds (with only pre-constant modified) as long as $d=O(\mathrm{poly}(k))$, which can be readily seen from the following inequalities that hold for any $\varepsilon>0$:
    $N^{\mathsf{box}\text{-}\mathsf{cover}}(\mathcal{X}_{\mathsf{data}},\|\cdot\|_{2},\varepsilon)$ $\leq N^{\mathsf{cover}}(\mathcal{X}_{\mathsf{data}},\|\cdot\|_{2},\varepsilon/2)\leq N^{\mathsf{box}\text{-}\mathsf{cover}}(\mathcal{X}_{\mathsf{data}},\|\cdot\|_{2},d^{-1/2}\varepsilon),$
    $N^{\mathsf{pack}}(\mathcal{X}_{\mathsf{data}},\|\cdot\|_{2},\varepsilon)$ $\leq N^{\mathsf{cover}}(\mathcal{X}_{\mathsf{data}},\|\cdot\|_{2},\varepsilon/2)\leq N^{\mathsf{pack}}(\mathcal{X}_{\mathsf{data}},\|\cdot\|_{2},\varepsilon/4).$
    In sum, our results remain valid when our intrinsic dimension is replaced by the Minkowski dimension.
  • Sparse vectors. The concept of sparsity is widely used in high-dimensional statistics to characterize low-dimensional structures. We say a vector $x\in\mathbb{R}^{d}$ is $s$-sparse if it has at most $s$ non-zero entries. Denoting $\mathcal{S}^{d-1}=\{x\in\mathbb{R}^{d}\mid\|x\|_{2}=1\}$ to be the unit Euclidean sphere in $\mathbb{R}^{d}$, it is known that sparse vectors on $\mathcal{S}^{d-1}$ has low metric entropy (Vershynin, 2009), which implies the following lemma.
    Lemma 4.
    Suppose $s<d/2$, and choose $k=s\log(d)$ in Assumption 1. If $\mathcal{X}_{\mathsf{data}}$ is contained within an $\varepsilon_{0}$-enlargement of the set of $s$-sparse vectors in $\mathcal{S}^{d-1}$, then Assumption 1 holds.
  • Union of the above sets. Our framework can also cover a union of the structures mentioned above, that is, Assumption 1 automatically holds when the data are supported on the union of $k^{O(k)}$ subsets, each of which satisfies Assumption 1. For the ImageNet dataset, the intrinsic dimension is estimated to be $k\approx 43$ by means of the Levina-Bickel method (Levina and Bickel, 2004), which shares similar flavor as the doubling dimension mentioned above. In this case, the number of subsets allowed in our framework could be as large as $10^{70}$, which manifests the flexibility and scalability of our analysis.
@@ 4.1.2 Discretization time points
Recall that the discretization time points are $0=t_{0}<t_{1}<\cdots<t_{N}<t_{N+1}=T$, with $T$ the time horizon of the forward process. We also define the early stopping time
@@
Following prior diffusion model theory (e.g., Benton et al. (2024)), we introduce a key quantity, $\kappa$, defined to be the smallest quantity obeying
@@
To get a sense of the size of $\kappa$, it is helpful to bear in mind the following schedule that has been adopted in recent theoretical analysis (Chen et al., 2023a; Benton et al., 2024; Li et al., 2025b):
$t_{n}$ $=\begin{cases}\frac{2(T-1)n}{N},\qquad&\text{ if }0\leq n\leq N/2,\\ T-\delta^{2(n-N/2)/N},&~\text{if }N/2<n\leq N,\end{cases}$ (15)
where we assume $N$ to be even for simplicity. In words, this schedule of time points can be divided into two phases: the first phase consists of the first $N/2$ time instances that are linearly spaced within [0,T1][0,T-1], whereas the second phase consists of the remaining $N/2$ time points that whose distance to $T$ decays exponentially from 1 to $\delta$. As shown in Benton et al. (2024, Corollary 1), this choice (15) satisfies
@@
For this reason, we can often interpret $\kappa$ as the average time interval, or simply $\widetilde{O}(1/N)$ as $T$ is typically taken to be $\widetilde{O}(1)$. In some other diffusion model literature, a uniform discretization schedule with rescaled forward and backward process is considered. Our analysis techniques can be easily adapted to this setting, and the following results, like Corollary 1, remain valid.
@@ 4.1.3 Score estimation errors
Assume the weighted sum of the mean squared score estimation errors is finite, i.e.,
@@
Treating the score learning phase as a black box, Assumption 3 singles out a single quantity, $\varepsilon_{\mathsf{score}}\geq 0$, that captures the accuracy of the pre-trained score functions, which is commonly used in recent development of diffusion model theory (Chen et al., 2023a; Li et al., 2024b, c; Benton et al., 2024; Dou et al., 2024; Li and Yan, 2024b; Li et al., 2024a; Li and Yan, 2024a; Potaptchik et al., 2024; Li and Jiao, 2024; Azangulov et al., 2024; Liang et al., 2024). Notably, assuming a weighted average of the mean squared score errors is clearly less stringent than requiring the mean squared score estimation errors to hold uniformly over all time discretization points (Chen et al., 2022b; Lee et al., 2022; Chen et al., 2023c; Wu et al., 2024).
@@ 4.2 Convergence guarantees for DDPM
  • $p_{\mathsf{out}}$: the distribution of $Y_{t_{N}}$ (or equivalently, the distribution of the DDPM output after $N$ steps).
@@
Suppose that Assumptions 1, 2 and 3 hold. Assume that $T>1$, $\kappa\leq 0.9$, $1/\mathrm{poly}(k)\lesssim\delta<1$ and $d=O\big(\mathrm{poly}(k)\big)$. Then the DDPM sampler (2) satisfies
$\mathsf{KL}(q_{\delta}\,||\,p_{\mathsf{out}})\lesssim\kappa\min\left\{\mathbb{E}\big[\|X_{0}\|_{2}^{2}\big],Tk\log k\right\}+\kappa^{2}Nk\log k+\varepsilon_{\mathsf{score}}^{2}+\big(d+\mathbb{E}\big[\|X_{0}\|_{2}^{2}\big]\big)e^{-2T}.$ (17)
The KL divergence upper bound in Theorem 2 can be divided into three parts. The first two terms on the right-hand side of (17) capture the time discretization error of the DDPM dynamics (namely, the error incurred when discretizing (12) as (11b)), the third term on the right-hand side of (17) reflects the effect of the score estimation error, whereas the fourth term corresponds to the initialization error (namely, the error induced by drawing $Y_{0}$ from $\mathcal{N}(0,I_{d})$ rather than the distribution of $X_{T}$). Note that we are comparing the DDPM output with $q_{\delta}$ (i.e., the distribution of a slightly noisy version of the data) rather than $p_{\mathsf{data}}=q_{0}$, as adding a little Gaussian noise helps avoid some highly irregular cases. For this reason, $\delta>0$ is commonly referred to as the early stopping time in prior theory. Note that $\delta$ can be taken to be very small, ensuring that $q_{\delta}$ is exceedingly close to $p_{\mathsf{data}}$. The assumption $\kappa\leq 0.9$ can be replaced by a more general condition $\kappa\leq 1-\varepsilon_{\kappa}$ for any small constant $\varepsilon_{\kappa}>0$, which still suffices for the analysis.
@@
$\mathsf{KL}(q_{\delta}\,||\,p_{\mathsf{out}})\lesssim\varepsilon^{2}+\varepsilon_{\mathsf{score}}^{2},$
provided that $1/\mathrm{poly}(k)\lesssim\delta<1$, $d=O\big(\mathrm{poly}(k)\big)$, and
@@
$N=\widetilde{\Theta}(k/\varepsilon^{2})~~\text{steps}$
@@
to approximate the distribution of $X_{\delta}$ (i.e., the target data distribution $p_{\mathsf{data}}$ convoluted with exceedingly small Gaussian noise) to within $\varepsilon^{2}$ KL divergence. Encouragingly, this iteration complexity is linear in the intrinsic dimension $k$ (modulo some logarithmic factor), which reduces to the state-of-the-art theory when $k=d$ (i.e., the case without low-dimensional structure) (Benton et al., 2024). It is noteworthy that all this is achieved without any prior knowledge about the potential low-dimensional structure in $p_{\mathsf{data}}$, thereby unraveling the remarkable capability of DDPM to adapt to unknown low dimensionality. It is also worth noting that this linear dependency in $k$ holds without any additional burn-in requirement; that is, its validity is guaranteed without requiring $N$ to first exceed, say, some polynomial function in $k$ or $d$.
@@
To understand the optimality of DDPM in adapting to low dimensionality, it is helpful to first revisit the lower bounds established by Chen et al. (2022b, Theorem 7) and Benton et al. (2024, Appendix H) for the full-dimensional case $k=d$; namely, there exists a simple $d$-dimensional distribution $p_{\mathsf{data}}$ for which the KL error is lower bounded by some function scaling linearly in $d$. These arguments can be straightforwardly extended to the low-dimensional case, revealing the existence of a distribution $p_{\mathsf{data}}$ with intrinsic dimension $k$, that requires an iteration complexity scaling at least linearly with $k$, regardless of the sampler in use.
Comparison with the concurrent work Potaptchik et al. (2024).
@@
Assuming the data lies on a low-dimensional manifold, this concurrent work, posted two weeks before ours, established an iteration complexity that scales linearly in the manifold dimension. As alluded to previously, our results and theirs share clearly similarities in both the aims and the results. Here, let us also point out several differences here. In Potaptchik et al. (2024), the data is assumed to be strictly supported on a constant-scale manifold with diameter less than $1$ (see Assumptions A and B therein), and the density function (w.r.t. to the standard volume measure) needs to be both upper and lower bounded uniformly (see Assumption C therein). In comparison, we consider a more general setting where the data support is only required to stay close to a low-dimensional structure whose size can grow polynomially in $k$. Also, no boundedness requirement is imposed on the density function in our theory, thereby accommodating a much broader scenario. Notably, the theory in Potaptchik et al. (2024) only applies to data with $\mathbb{E}\big[\|X_{0}\|_{2}^{2}\big]=O(1)$ second moment (as a consequence of their assumption that the diameter of the manifold is $O(1)$), while ours is able to accommodating polynomia lly large second moment $\mathbb{E}\big[\|X_{0}\|_{2}^{2}\big]=\mathrm{poly}(k)$. Our proof invokes the covering-style argument to accommodate broader low-dimensional structure, which differs from the techniques in Potaptchik et al. (2024).